### v1

@@ 6 Numerical experiments
In this section, we validate the effectiveness of the DCEN method through numerical experiments on sparse signal recovery, MRI image reconstruction, and high-dimensional variable selection. All computational experiments were carried out on a PC equipped with an Intel Core i9–12900H processor (2.50 GHz), and the simulation environment was implemented on the MATLAB R2025b platform. We compare the DCEN model with the state-of-the-art models and algorithms in sparse recovery, including $\ell_{1}$-ADMM, $\ell_{p}(p=1/2)$, IRLS-$\ell_{p}(p=1/2)$, $\ell_{1}/\ell_{\infty}$-FISTA, $\ell_{1}/\ell_{\infty}$-ADMM, $\ell_{1}-\alpha\ell_{2}$-ADMM, $\ell_{1}-\alpha\ell_{2}$-IAADMM ($\alpha$ is iteratively adjusted), $\ell_{1}-\alpha\ell_{2}$-DCA, and $\ell_{1}-\alpha\ell_{2}$-FBS, for noiseless and noisy cases in subsections 6.2 and 6.3, respectively.
6.1 Parameter settings
@@
  • Oversampled Discrete Cosine Transform (DCT) Matrix: Following the construction methods in [rahimi2019scale, tao2022minimization, wang2023variant, yin2015minimization], the column vectors $\bm{a}_{j}$ are defined as $\bm{a}_{j}:=1/\sqrt{m}\cdot\cos\left(2\pi\bm{w}j/F\right),\;j=1,\dots,n$, where $\bm{w}\in[0,1]^{m}$ is a random vector uniformly distributed within the unit hypercube. The parameter $F>0$ controls the coherence of the matrix; a larger value of $F$ implies a higher correlation between column vectors, thereby increasing the difficulty of the recovery problem.
  • Gaussian Matrix: This class of matrices is generated by sampling from a multivariate normal distribution $\mathcal{N}(0,\bm{\Sigma})$. By the settings in [yin2015minimization], the elements of the covariance matrix $\bm{\Sigma}$ are given by $\Sigma_{i,j}=\{(1-r)*\bm{I}(i=j)+r\}_{i,j}$, where $r\in(0,1)$ is the correlation coefficient. A higher value of $r$ indicates strong correlation among observations, which is typically regarded as a challenging scenario for sparse recovery.
We set $\epsilon=10^{-6}$ and the maximum number of iterations $K=50$ for DCEN-DCA in Algorithm 1. For the DCEN-DCA subproblem in Algorithm 2 and Algorithm 3, we chose parameters $\varepsilon_{\mathrm{abs}}=10^{-6}$, $\varepsilon_{\mathrm{rel}}=10^{-6}$, $\varepsilon=10^{-6}$, and the maximum number of iterations $T=5n$. The parameter settings for DCA and ADMM, utilized for solving the $\ell_{1}-\alpha\ell_{2}$ model, are maintained consistent with those employed in DCEN. The same goes for $\ell_{1}/\ell_{\infty}$-ADMM. For $\ell_{1}/\ell_{\infty}$-FISTA, the algorithm is terminated once tolerance $\text{tol}=10^{-6}$ is satisfied or when the number of iterations reaches $5n$. All other settings of the algorithm were set to default ones. To obtain a high-quality initial estimate and avoid local minima, we adopt a warm-start strategy. Specifically, following the standard practice adopted in existing literature, we initialize all algorithms using the solution of $\ell_{1}$-ADMM [rahimi2019scale, tao2022minimization, yin2015minimization]. a We assess sparse recovery performance using the success rate, defined as the ratio of successful reconstructions to the total number of trials. A trial is considered successful when the relative error between the reconstruction vector $\bm{x}^{*}$ and the ground truth signal $\bm{x}^{\sharp}$ satisfies $\frac{\|\bm{x}^{*}-\bm{x}^{\sharp}\|_{2}}{\|\bm{x}^{\sharp}\|_{2}}<10^{-3}$.
6.2 Efficiency Evaluation in Noiseless Scenario
This subsection is dedicated to comparing the efficiency of the DCEN model and its corresponding solver against established methods in a noiseless setting. The regularization parameters for all compared models are uniformly fixed at $\lambda=10^{-7}$, with the exception of the $\ell_{p}$ model, which employs an adaptive parameter selection strategy. Apart from the specific stopping criteria mentioned in Section 6.1, all remaining algorithmic parameters adopt their respective default settings.
@@ 6.2 Efficiency Evaluation in Noiseless Scenario
To enable a more systematic comparison, we unify the matrix dimensions to $64\times 1024$ for both the oversampled DCT and Gaussian random matrices. The oversampling factor $F$ is set to 20 and 30, while the correlation parameter $r$ for the Gaussian random matrix is set to 0.2 and 0.9. The sparsity level $s$ is varied from 2 to 30 in increments of 2. Each configuration is again repeated 100 times. As shown in Figure 3(a), DCEN-DCA achieves significantly higher success rates than the $\ell_{1}$-$\alpha\ell_{2}$ model and other competing algorithms in high coherence scenarios. For Gaussian random matrices, the IRLSLp model exhibits excellent performance, consistent with the observations in [tao2022minimization] and [wang2020accelerated]. Notably, both the IRLSLp and $\ell_{1/2}$ models—solved via iteratively reweighted algorithms and half-thresholding, respectively—require prior knowledge of the true sparsity level $s$ of the signal $\bm{x}^{\sharp}$. In our experiments, this ground-truth sparsity is provided. Leveraging this crucial prior information, the IRLSLp model attains substantially higher success rates compared to methods that do not assume sparsity knowledge.
(a) Oversampled DCT matrix: $F=20$ and sparsity $s=20$.
(b) Gaussian matrix: $r=0.2$ and sparsity $s=10$.
(c) Oversampled DCT matrix: $F=30$ and sparsity $s=10$.
(d) Gaussian matrix: $r=0.9$ and sparsity $s=15$.
@@
(a) Oversampled DCT matrix.
(b) Gaussian matrix.
Figure 3: Success rate versus sparsity $s$ for oversampled DCT and Gaussian matrices.
6.3 Robust recovery of DCEN
@@
Table 1: Reconstruction SNR (dB) for the DCEN and $\ell_{1}-\alpha\ell_{2}$ models and their solvers under varying noise levels using an oversampled DCT sensing matrix.
@@
Table 2: Reconstruction SNR (dB) for the DCEN and $\ell_{1}-\alpha\ell_{2}$ models and their solvers under varying noise levels using a high-coherence Gaussian random sensing matrix ($r=0.9$).
@@
Experiments are conducted on two types of sensing matrices. For the oversampled DCT matrix, we set its dimensions to $100\times 1000$, with an oversampling factor $F=15$, sparsity level $s=10$, and a minimum separation of 15 between non-zero entries. For the Gaussian random matrix, the dimensions are also $100\times 1000$, with a correlation parameter $r=0.9$ and sparsity level $s=30$. Under each experimental configuration, we perform 50 independent trials and report the average reconstruction SNR across all runs. Table 1 presents the results for the oversampled DCT matrix. The regularization parameters $\lambda$ and weight parameter $\gamma$ are tuned according to the noise level using a grid search strategy. The results demonstrate that, in most cases, particularly when solved via DCA, DCEN achieves higher reconstruction SNR compared to the $\ell_{1}$-$\alpha\ell_{2}$ model. Table 2 reports the results for the Gaussian random matrix. Under high-coherence setting ($r=0.9$), DCEN-DCA exhibits a clear advantage, consistently outperforming both the $\ell_{1}$-$\alpha\ell_{2}$ model and ADMM-based variants. Notably, DCEN-DCA not only yields higher reconstruction accuracy but also demonstrates superior robustness across varying noise levels. In contrast, the performance of ADMM-type algorithms degrades, likely due to the adverse effects of high matrix coherence on convergence and algorithm parameters tuning. Overall, across all tested scenarios, DCEN-DCA consistently outperforms both ADMM-based solvers and the standard $\ell_{1}$-$\alpha\ell_{2}$ model.
6.4 MRI Image Reconstruction
@@
(47) $\min_{\bm{u}}\;\gamma\bigl(\|\nabla\bm{u}\|_{1}-\alpha\|\nabla\bm{u}\|_{2}\bigr)+(1-\gamma)\|\nabla\bm{u}\|_{2}^{2}\quad\text{s.t.}\quad\bm{R}\mathcal{F}\bm{u}=\bm{f},$
@@
(48) $\min_{\bm{u}}\;\gamma\bigl(|\partial_{\bm{x}}\bm{u}|+|\partial_{\bm{y}}\bm{u}|\bigr)+(1-\gamma)\bigl(|\partial_{\bm{x}}\bm{u}|^{2}+|\partial_{\bm{y}}\bm{u}|^{2}\bigr)-\alpha\gamma\langle\bm{t}^{k},\nabla\bm{u}\rangle\;\text{s.t.}\;\bm{R}\mathcal{F}\bm{u}=\bm{f},$
where the gradient direction vector $\bm{t}^{k}=(\bm{t}_{\bm{x}^{k}},\bm{t}_{\bm{y}^{k}})$ is given by
(49) $\bm{t}^{k}=(\partial_{\bm{x}}\bm{u}^{k},\partial_{\bm{y}}\bm{u}^{k})/\sqrt{|\partial_{\bm{x}}\bm{u}^{k}|^{2}+|\partial_{\bm{y}}\bm{u}^{k}|^{2}}.$
To solve subproblem (48) efficiently, we introduce auxiliary variables $\bm{d}_{\bm{x}}=\partial_{\bm{x}}\bm{u}$ and $\bm{d}_{\bm{y}}=\partial_{\bm{y}}\bm{u}$, and reformulate the problem (48) via quadratic penalization as
(50) $(\bm{u}_{k+1},$ $\bm{d}_{\bm{x}_{k+1}},\bm{d}_{\bm{y}_{k+1}})$
$=\arg\min_{\bm{u},\bm{d_{x}},\bm{d_{y}}}\;\gamma(|\bm{d_{x}}|+|\bm{d_{y}}|)+(1-\gamma)(\bm{d_{x}}^{2}+\bm{d_{y}}^{2})-\gamma\alpha(\bm{t_{x}}^{k}\bm{d_{x}}+\bm{t_{y}}^{k}\bm{d_{y}})$
@@
By introducing Bregman multipliers $\bm{b}=(\bm{b_{x}},\bm{b_{y}})$ for the constraint $\bm{d}=\nabla\bm{u}$ and a dual variable $\bm{z}$ for the data consistency constraint $\bm{R}\mathcal{F}\bm{u}=\bm{f}$, the split Bregman framework can be applied to solve (50) efficiently. The complete algorithmic procedure is summarized in Algorithm 4. Notably, when $\gamma=1$, the model (47) reduces to the $\ell_{1}-\alpha\ell_{2}$ TV-type formulation, and the proposed algorithm simplifies to Algorithm 3 presented in [yin2015minimization].
@@
5:   $\bm{u}\leftarrow\left(\mu\bm{R}^{\top}\bm{R}-\beta\mathcal{F}\bm{\Delta}\mathcal{F}^{\top}\right)^{-1}\left(\mu\mathcal{F}^{\top}\bm{R}\bm{z}+\beta\bm{D_{x}}^{\top}(\bm{d_{x}}-\bm{b_{x}})+\beta\bm{D_{y}}^{\top}(\bm{d_{y}}-\bm{b_{y}})\right)$
6:   Compute $\bm{d_{x}}\leftarrow\mathrm{shrink}\!\left(\dfrac{\gamma\alpha\bm{t_{x}}+\beta(\bm{D_{x}u}+\bm{b_{x}})}{\beta+2(1-\gamma)},\,\dfrac{\gamma}{\beta+2(1-\gamma)}\right)$
7:   Compute $\bm{d_{y}}\leftarrow\mathrm{shrink}\!\left(\dfrac{\gamma\alpha\bm{t_{y}}+\beta(\bm{D_{y}u}+\bm{b_{y}})}{\beta+2(1-\gamma)},\,\dfrac{\gamma}{\beta+2(1-\gamma)}\right)$
8:   Update $\bm{b_{x}}\leftarrow\bm{b_{x}}+\bm{D_{x}u}-\bm{d_{x}}$
9:   Update $\bm{b_{y}}\leftarrow\bm{b_{y}}+\bm{D_{y}u}-\bm{d_{y}}$
10:  end for
11:  Update $\bm{z}\leftarrow\bm{z}+\bm{f}-\bm{R}\mathcal{F}\bm{u}$
12:end for
@@
We employ the standard Shepp-Logan phantom as the test image for our experiments. The comparative models include the classical TV-$\ell_{1}$[candes2006robust], MCTV-$\ell_{1}$ [liu2018convex], TV-$\ell_{2}$[shen2021magnetic], MCTV-$\ell_{2}$[shen2021magnetic], $\ell_{1}$-$\ell_{2}$, and $\ell_{1}/\ell_{2}$ models. Figure 4 presents the reconstruction accuracy of all models under a radial sampling scheme with only eight projection lines. Experimental results demonstrate that DCEN-TV achieves the highest reconstruction accuracy. Compared to the $\ell_{1}$-$\ell_{2}$ model, the relative error of DCEN is reduced by an order of magnitude, indicating a significant improvement in reconstruction quality. Notably, although the $\ell_{1}/\ell_{2}$ model is currently regarded as state-of-the-art in MRI reconstruction, it yields substantially higher relative errors than DCEN—highlighting the superior performance of our method in sparse MRI reconstruction.
6.5 Variable Selection Performance under Highly Correlated Predictors
To comprehensively evaluate the feature selection and sparse recovery performance of various sparse regression models in highly correlated scenarios, we conduct the following numerical experiment. Specifically, the sample size is set to $n=20$ and the total number of variables is $p=100$. The first three predictors are generated from a multivariate normal distribution $\mathcal{N}(0,\Sigma)$ with covariance matrix $\Sigma$ having off-diagonal entries $\rho=0.99$, forming a highly correlated signal block. The remaining 97 predictors are i.i.d. standard normal noise. The true regression coefficients are nonzero only for the first three variables. Gaussian noise with standard deviation 1.2 is added to the response. All results are averaged over 100 independent Monte Carlo trials. We report the selection probabilities of the three true variables (Var1 to Var3), the average total number of selected variables, the average number of falsely selected noise variables, and the average selection rate of noise variables. The results are summarized in Table 3. Figure 5 reports the results obtained from 100 independent Monte Carlo simulations.
From Table 3, we know that LASSO and the $\ell_{1}-\alpha\ell_{2}$ penalty suffer from severely degraded selection probabilities under strong correlation (average true-variable recovery rates of only 29.3% and 24.3%, respectively). Elastic Net substantially improves the balance across correlated variables (average 76.0%), but at the cost of a markedly higher false discovery rate (8.65 noise variables on average). Several nonconvex penalties (e.g., $\ell_{1}/\ell_{\infty}$ via FISTA and $\ell_{1}/\ell_{2}$) exhibit extreme imbalance among correlated predictors, often concentrating nearly all nonzero mass on a single variable while discarding the others. DCEN-ADMM achieves the best overall trade-off: it attains true-variable selection probabilities of 85.0%, 78.0%, and 86.0% (average 83.0%, significantly superior to Elastic Net), while keeping the average number of falsely selected noise variables at only 4.98 and the noise selection rate at 5.13%. Compared with the DCA-based implementation of DCEN, ADMM demonstrates considerably greater stability in this highly challenging correlated regime. These findings indicate that DCEN-ADMM effectively mitigates the instability of traditional convex approaches and the tendency of purely nonconvex penalties to converge to poor local solutions. Consequently, it delivers substantially more reliable support recovery in the small-sample, high-dimensional, strongly correlated setting—a scenario of particular practical relevance in bioinformatics and imaging applications.

### v2

@@ VII Numerical experiments
In this section, we validate the effectiveness of the DCEN method through numerical experiments on sparse signal recovery, MRI image reconstruction, and high-dimensional variable selection. All computational experiments were carried out on a PC equipped with an Intel Core i9–12900H processor (2.50 GHz), and the simulation environment was implemented on the MATLAB R2025b platform. We compare the DCEN model with the state-of-the-art models and algorithms in sparse recovery, including $\ell_{1}$-ADMM, $\ell_{p}(p=1/2)$, IRLSLp(p=1/2)(p=1/2), $\ell_{1}/\ell_{\infty}$-FISTA, $\ell_{1}/\ell_{\infty}$-ADMM, $\ell_{1}-\alpha\ell_{2}$-ADMM, $\ell_{1}-\alpha\ell_{2}$-IAADMM ($\alpha$ is iteratively adjusted), and $\ell_{1}-\alpha\ell_{2}$-DCA, for noiseless and noisy cases in subsections VII-B and VII-C, respectively.
VII-A Parameter settings
@@
  • Oversampled Discrete Cosine Transform (DCT) Matrix: Following the construction methods in [28, 31, 36, 41], the column vectors $\bm{a}_{j}$ are defined as $\bm{a}_{j}:=1/\sqrt{m}\cdot\cos\left(2\pi\bm{w}j/F\right),\;j=1,\dots,n$, where $\bm{w}\in[0,1]^{m}$ is a random vector uniformly distributed within the unit hypercube. The parameter $F>0$ controls the coherence of the matrix; a larger value of $F$ implies a higher correlation between column vectors, thereby increasing the difficulty of the recovery problem.
  • Gaussian Matrix: This class of matrices is generated by sampling from a multivariate normal distribution $\mathcal{N}(0,\bm{\Sigma})$. By the settings in [41], the elements of the covariance matrix $\bm{\Sigma}$ are given by $\Sigma_{i,j}=\{(1-r)*\bm{I}(i=j)+r\}_{i,j}$, where $r\in(0,1)$ is the correlation coefficient. A higher value of $r$ indicates strong correlation among observations, which is typically regarded as a challenging scenario for sparse recovery.
We set $\epsilon=10^{-6}$ and the maximum number of iterations $K=50$ for DCEN-DCA in Algorithm 1. For the DCEN-DCA subproblem in Algorithm 2 and Algorithm 3, we chose parameters $\varepsilon_{\mathrm{abs}}=10^{-6}$, $\varepsilon_{\mathrm{rel}}=10^{-6}$, $\varepsilon=10^{-6}$, and the maximum number of iterations $T=5n$. The parameter settings for DCA and ADMM, utilized for solving the $\ell_{1}-\alpha\ell_{2}$ model, are maintained consistent with those employed in DCEN. The same goes for $\ell_{1}/\ell_{\infty}$-ADMM. For $\ell_{1}/\ell_{\infty}$-FISTA, the algorithm is terminated once tolerance $\text{tol}=10^{-6}$ is satisfied or when the number of iterations reaches $5n$. All other settings of the algorithm were set to default ones. To obtain a high-quality initial estimate and avoid local minima, we adopt a warm-start strategy. Specifically, following the standard practice adopted in existing literature, we initialize all algorithms using the solution of $\ell_{1}$-ADMM [28, 31, 41]. We assess sparse recovery performance using the success rate, defined as the ratio of successful reconstructions to the total number of trials. A trial is considered successful when the relative error between the reconstruction vector $\bm{x}^{*}$ and the ground truth signal $\bm{x}^{\sharp}$ satisfies $\frac{\|\bm{x}^{*}-\bm{x}^{\sharp}\|_{2}}{\|\bm{x}^{\sharp}\|_{2}}<10^{-3}$.
VII-B Efficiency Evaluation in Noiseless Scenario
This subsection is dedicated to comparing the efficiency of the DCEN model and its corresponding solver against established methods in a noiseless setting. The regularization parameters for all compared models are uniformly fixed at $\lambda=10^{-7}$, with the exception of the $\ell_{p}$ model, which employs an adaptive parameter selection strategy. Apart from the specific stopping criteria mentioned in Section VII-A, all remaining algorithmic parameters adopt their respective default settings.
@@ VII-B Efficiency Evaluation in Noiseless Scenario
To enable a more systematic comparison, we unify the matrix dimensions to $64\times 1024$ for both the oversampled DCT and Gaussian random matrices. The oversampling factor $F$ is set to 20 and 30. The sparsity level $s$ is varied from 2 to 30 in increments of 2. Each configuration is again repeated 100 times. As shown in Figure 2, DCEN-DCA achieves higher success rates than the $\ell_{1}-\alpha\ell_{2}$ model and other competing algorithms in high coherence scenarios.
@@
(a) Oversampled DCT: $F=20$ and $s=20$.
(b) Gaussian: $r=0.2$ and $s=10$.
(c) Oversampled DCT: $F=30$ and $s=10$.
(d) Gaussian: $r=0.9$ and $s=15$.
Figure 2: Success rate versus sparsity $s$ for oversampled DCT.
VII-C Robust recovery of DCEN
@@
TABLE I: Reconstruction SNR (dB) for the DCEN and $\ell_{1}-\alpha\ell_{2}$ models and their solvers under varying noise levels using an oversampled DCT sensing matrix.
@@
TABLE II: Reconstruction SNR (dB) for the DCEN and $\ell_{1}-\alpha\ell_{2}$ models and their solvers under varying noise levels using a high-coherence Gaussian random sensing matrix ($r=0.9$).
@@
Experiments are conducted on two types of sensing matrices. For the oversampled DCT matrix, we set its dimensions to $100\times 1000$, with an oversampling factor $F=15$, sparsity level $s=10$, and a minimum separation of 15 between non-zero entries. For the Gaussian random matrix, the dimensions are also $100\times 1000$, with a correlation parameter $r=0.9$ and sparsity level $s=30$. Under each experimental configuration, we perform 50 independent trials and report the average reconstruction SNR across all runs. Table I presents the results for the oversampled DCT matrix. The regularization parameters $\lambda$ and weight parameter $\gamma$ are tuned according to the noise level using a grid search strategy. The results demonstrate that, in most cases, particularly when solved via DCA, DCEN achieves higher reconstruction SNR compared to the $\ell_{1}-\alpha\ell_{2}$ model. Table II reports the results for the Gaussian random matrix. Under high-coherence setting ($r=0.9$), DCEN-DCA exhibits a clear advantage, consistently outperforming both the $\ell_{1}-\alpha\ell_{2}$ model and ADMM-based variants. Notably, DCEN-DCA not only yields higher reconstruction accuracy but also demonstrates superior robustness across varying noise levels. In contrast, the performance of ADMM-type algorithms degrades, likely due to the adverse effects of high matrix coherence on convergence and algorithm parameters tuning. Overall, across all tested scenarios, DCEN-DCA consistently outperforms both ADMM-based solvers and the standard $\ell_{1}-\alpha\ell_{2}$ model.
VII-D MRI Image Reconstruction
@@
$\min_{\bm{u}}\;\gamma\bigl(\|\nabla\bm{u}\|_{1}-\alpha\|\nabla\bm{u}\|_{2}\bigr)+(1-\gamma)\|\nabla\bm{u}\|_{2}^{2}\quad\text{s.t.}\quad\bm{R}\mathcal{F}\bm{u}=\bm{f},$ (53)
@@
$\min_{\bm{u}}\;r_{1}(\bm{u})+r_{2}(\bm{u})-\alpha\gamma\langle\bm{t}^{k},\nabla\bm{u}\rangle\;\text{s.t.}\;\bm{R}\mathcal{F}\bm{u}=\bm{f},$ (54)
where $r_{1}(\bm{u})=\gamma\bigl(|\partial_{\bm{x}}\bm{u}|+|\partial_{\bm{y}}\bm{u}|\bigr)$, $r_{2}(\bm{u})=(1-\gamma)\bigl(|\partial_{\bm{x}}\bm{u}|^{2}+|\partial_{\bm{y}}\bm{u}|^{2}\bigr)$ and the gradient direction vector $\bm{t}^{k}=(\bm{t}_{\bm{x}^{k}},\bm{t}_{\bm{y}^{k}})$ is given by
$\bm{t}^{k}=(\partial_{\bm{x}}\bm{u}^{k},\partial_{\bm{y}}\bm{u}^{k})/\sqrt{|\partial_{\bm{x}}\bm{u}^{k}|^{2}+|\partial_{\bm{y}}\bm{u}^{k}|^{2}}.$ (55)
To solve subproblem (54) efficiently, we introduce auxiliary variables $\bm{d}_{\bm{x}}=\partial_{\bm{x}}\bm{u}$, $\bm{d}_{\bm{y}}=\partial_{\bm{y}}\bm{u}$, $\bm{\bar{d}}=(\bm{d}_{\bm{x}},\bm{d}_{\bm{y}})^{\top}$ and reformulate the problem (54) via quadratic penalization as
$(\bm{u}_{k+1},\bm{d}_{\bm{x}_{k+1}},\bm{d}_{\bm{y}_{k+1}})$ (56)
$=\arg\min_{\bm{u},\bm{d_{x}},\bm{d_{y}}}\;\gamma r_{1}(\bm{\bar{d}})+r_{2}(\bm{\bar{d}})-\gamma\alpha(\bm{t_{x}}^{k}\bm{d_{x}}+\bm{t_{y}}^{k}\bm{d_{y}})$
@@
By introducing Bregman multipliers $\bm{b}=(\bm{b_{x}},\bm{b_{y}})$ for the constraint $\bm{d}=\nabla\bm{u}$ and a dual variable $\bm{z}$ for the data consistency constraint $\bm{R}\mathcal{F}\bm{u}=\bm{f}$, the split Bregman framework can be applied to solve (56) efficiently. The complete algorithmic procedure is summarized in Algorithm 4. Notably, when $\gamma=1$, the model (53) reduces to the $\ell_{1}-\alpha\ell_{2}$ TV-type formulation, and the proposed algorithm simplifies to Algorithm 3 presented in [41].
@@
5:   $\bm{M}_{R}=\mu\mathcal{F}^{\top}\bm{R}\bm{z}+\beta\bm{D_{x}}^{\top}(\bm{d_{x}}-\bm{b_{x}})+\beta\bm{D_{y}}^{\top}(\bm{d_{y}}-\bm{b_{y}})$
6:   $\bm{u}\leftarrow(\mu\bm{R}^{\top}\bm{R}-\beta\mathcal{F}\bm{\Delta}\mathcal{F}^{\top})^{-1}\bm{M}_{R}$
7:   $\bm{d_{x}}\leftarrow\mathrm{shrink}(\dfrac{\gamma\alpha\bm{t_{x}}+\beta(\bm{D_{x}u}+\bm{b_{x}})}{\beta+2(1-\gamma)},\,\dfrac{\gamma}{\beta+2(1-\gamma)})$
8:   $\bm{d_{y}}\leftarrow\mathrm{shrink}(\dfrac{\gamma\alpha\bm{t_{y}}+\beta(\bm{D_{y}u}+\bm{b_{y}})}{\beta+2(1-\gamma)},\,\dfrac{\gamma}{\beta+2(1-\gamma)})$
9:   Update $\bm{b_{x}}\leftarrow\bm{b_{x}}+\bm{D_{x}u}-\bm{d_{x}}$
10:   Update $\bm{b_{y}}\leftarrow\bm{b_{y}}+\bm{D_{y}u}-\bm{d_{y}}$
11:  end for
12:  Update $\bm{z}\leftarrow\bm{z}+\bm{f}-\bm{R}\mathcal{F}\bm{u}$
13:end for
@@
We employ the standard Shepp-Logan phantom as the test image for our experiments. The comparative models include the classical TV-$\ell_{1}$[4], MCTV-$\ell_{1}$ [24], TV-$\ell_{2}$[30], MCTV-$\ell_{2}$[30], $\ell_{1}$-$\ell_{2}$, and $\ell_{1}/\ell_{2}$ models. Figure 3 presents the reconstruction accuracy of all models under a radial sampling scheme with only eight projection lines. Experimental results demonstrate that DCEN-TV achieves the highest reconstruction accuracy. Compared to the $\ell_{1}$-$\ell_{2}$ model, the relative error of DCEN is reduced by an order of magnitude, indicating a significant improvement in reconstruction quality. Notably, although the $\ell_{1}/\ell_{2}$ model is currently regarded as state-of-the-art in MRI reconstruction, it yields substantially higher relative errors than DCEN—highlighting the superior performance of our method in sparse MRI reconstruction.
VII-E Variable Selection under Highly Correlated Predictors
To evaluate the feature selection and sparse recovery performance under high correlation, we conduct numerical experiments with $n=20$ and $p=100$. The simulation design incorporates a highly correlated signal block consisting of the first three predictors ($\rho=0.99$), while the remaining 97 variables are i.i.d. standard normal noise. Gaussian noise with $\sigma=1.2$ is added to the response, and all reported metrics are averaged over 100 independent Monte Carlo trials.
The experimental results, summarized in Table III and Figure 4, reveal distinct performance gaps among the tested models. Traditional methods, including LASSO and the $\ell_{1}-\alpha\ell_{2}$ penalty, exhibit severe degradation in selection probabilities, yielding average true-variable recovery rates of only 29.3% and 24.3%, respectively. While the Elastic Net improves the balance across correlated predictors with an average recovery rate of 76.0%, it does so at the expense of a markedly high false discovery rate, selecting an average of 8.65 noise variables. Furthermore, purely nonconvex penalties such as $\ell_{1}/\ell_{\infty}$ and $\ell_{1}/\ell_{2}$ demonstrate extreme instability, often concentrating nonzero mass on a single predictor while erroneously discarding the others.
In contrast, the proposed DCEN-ADMM achieves a superior trade-off between sensitivity and specificity. It attains individual true-variable selection probabilities of 85.0%, 78.0%, and 86.0% (averaging 83.0%), significantly outperforming the Elastic Net. Simultaneously, it maintains a low false discovery rate, with an average of only 4.98 falsely selected variables and a noise selection rate of 5.13%. Notably, the ADMM-based implementation exhibits greater stability than the DCA-based version in this challenging regime. These results suggest that DCEN-ADMM effectively mitigates the instability inherent in convex models and the local-minima issues of nonconvex approaches, providing robust support recovery for the small-sample, high-dimensional, and strongly correlated data common in bioinformatics and imaging applications.